# Towards a bottom-up formulation of spin kinetic theory

### v1

@@ 1 Introduction
In response, significant developments have been made in formulating the SKT for relativistic plasma Hattori:2019ahi ; Weickgenannt:2019dks ; Weickgenannt:2020aaf ; Weickgenannt:2021cuo ; Hayata2020 ; Sheng:2021kfc ; Yang:2020hri (see ref. Hidaka:2022dmn for a review). A key predecessor to SKT is chiral kinetic theory, which describes the distribution of axial density/helicity Son:2012wh ; Stephanov:2012ki ; Son:2012zy ; Chen:2012ca ; Chen:2013iga ; Manuel:2013zaa ; Manuel:2014dza ; Chen:2014cla ; Chen:2015gta ; Hidaka:2016yjf ; Huang:2018wdl . Most of the references mentioned above, particularly those from recent years, derive SKT using what we shall refer to as the ”top-down” approach. This approach begins with the microscopic Schwinger-Dyson equation for the Wigner function of the Fermionic field and then extracts the relevant equations for its components.
@@ 2.2.1 Dynamical variables
Because of the first inequality in eq. (2), we shall concentrate on the collisionless regime and neglect the self-energy corrections, which are of the order of $m_{D}$. The second inequality ensures that a gradient expansion exists in powers of $q/T_{{\rm eff}}$ (or $q/m_{{\rm eff}}$). Eq. (2) assumes the separation of scales: $m_{D},\gamma_{R}\ll T_{{\rm eff}}$, which is true for a weakly coupled plasma with coupling constant $g_{C}<1$ where $m_{D}\sim g_{C}T_{{\rm eff}}$ and $\gamma_{R}\sim g^{2}_{C}T_{\rm eff}$.
@@ 2.2.2 Poisson bracket and equation of motion
We assume that $\bm{g}$} has no background value, and consequently, neither does $\bm{\zeta}$}. In the absence of $\bm{\zeta}$}, the evolution of $\bm{g}$} is determined by $\epsilon$ in the same way as $f$. The presence of $\bm{\zeta}$ introduces the coupling between $\bm{g}$} and $f$ in their evolution. The term $\bm{g}\times\bm{\zeta}$ originates from the last term in eq. (6b) which is a manifestation of $SU(2)$ algebra of spin $1/2$ particle. This non-linear term implies that if $\bm{\zeta}\propto\bm{B}$, $\bm{g}$ will evolve unless it aligns with the magnetic field $\bm{B}$.
Summarizing, we adopt a ”bottom-up” viewpoint by proposing that dynamical variables of SKT, $(f,\gamma)$, obey E.o.Ms which are expressible in terms of spin-dependent P.B. Those equations have to be supplemented by the expression for the particle/spin energy $(\epsilon,{\bm{\zeta}})$. In general, $(\delta\epsilon,{\bm{\zeta}})$ are functionals of $(f,{\bm{g}})$ and external fields, similar to the framework of Landau Fermi liquid theory. The specification of those dependencies will be fixed by matching to the microscopic calculations; further details can be found in sec. 5.
@@ 2.2.3 Schwinger-Keldysh action
While one might think that $\Phi_{a}$ merely acts as the Lagrangian multiplier that ensures only solutions to E.o.M. contribute to the path integral  [^1] If we were using Martin–Siggia–Rose (MSR) method Martin:1973zz , we would place $i\phi_{a}$ where $\phi_{a}$ appears. However, the second condition in (9) requires that we shall use $\Phi_{a}$ without the prefactor $i$. , we emphasize that both $\Phi_{a}$ and its r-variable counterpart $\Phi$ possess much richer physical content in this formulation This is reflected through the fact that various symmetry transformations can be implemented by the linear shift of $\Phi_{r/a}$. Particularly, we are interested in $U(1)$ gauge transformation $A_{\mu}\to A_{\mu}+\partial_{\mu}\alpha(x)$ and space-time translation parametrized by $b_{\mu}$. By construction, they are given by (see also Ref Delacretaz:2025ifh )
@@ 2.4.2 Constitutive relation for axial Wigner function
We have projected $\bm{g}$} in the direction longitudinal and transverse to the three velocity $\bm{v}$, i.e., ${\bm{g}}_{\parallel}\equiv\hat{v}(\bm{v}\cdot{\bm{g}}),{\bm{g}}_{\perp}\equiv{\bm{g}}-{\bm{g}}_{\parallel}$. The coefficients $c_{0,1,2,3}$ are function of $\epsilon_{\bm{p}}$. We shall match eq. (22) with the field theoretical calculations later.
@@ 3 Linear response from kinetic theory
We shall match the kinetic theory description of the medium’s response to external disturbance with the field theory. In kinetic theory, the response is characterized by the deviations from the background distribution, i.e. $(\delta f,{\bm{g}})$ where $f=f_{0}+\delta f$ and $\bm{g}$}. Likewise, the presence of external fields also induces the change in energy, i.e., $\delta\epsilon$ in $\epsilon=\epsilon_{\bm{p}}+\delta\epsilon$ and $\bm{\zeta}$}. In our set-up, ${\bm{g}},\zeta$ has no background value so we do not put $\delta$ in those symbols for brevity.
To analyze linear response, we set up power counting to keep track of the external perturbation by introducing a small parameter $\delta$. The background value of the distribution function $f_{0},{\cal V}^{\mu}_{0}\sim{\cal O}(\delta^{0})$ while $(\delta\epsilon,\bm{\zeta},{\bm{g}})\sim{\cal O}(\delta)$. For consistency, the function that parametrizes the symmetry transformation is also ${\cal O}(\delta)$. While both $\delta$ and gradient are counted as small, they are independent. For example, among terms which are linear in $\delta$, $\delta f,\delta\epsilon\sim{\cal O}(\partial^{0})$ while $\zeta,g\sim{\cal O}(\partial)$.
@@
$\bm{g}(Q)$ $=\frac{\bm{v}\cdot\bm{q}}{v\cdot Q}\,f_{0}^{\prime}\,\bm{\zeta}(Q)\,.$ (25b)
We note from eqs. (25a),(25b) that $(\delta f,g)$ are non-analytic function in frequency and wave-vector $\omega,\bm{q}$, the point we shall return to in sec. 5. Note also that $(\delta f,{\bm{g}})$ in eq. (25) are proportional to $f^{\prime}_{0}$, which in the low temperature limit only receives contributions from excitations in the vicinity of the Fermi surface. The general lesson here is that the present EFT describes hard particles that contribute most to the deviation from the background value. The phase-space volume occupied by those excitations can be significantly smaller than that occupied by the particles in the bulk, indicating a reduction in the density of d.o.f.s governing the system’s dynamics.
@@ 4.1 Set-up
and $\sigma^{ab}=\frac{i}{2}[\gamma^{a},\gamma^{b}]_{-}$ are the generator of Lorentz transformation. The spin-connection term ensures the covariance of $W(x,{\cal P})$ under local Lorentz transformations (LLTs). The Wigner (Fourier) transform has been performed so that $W$ depends on both $x$ and the off-shell four-momentum ${\cal P}_{\mu}=(p_{0},\bm{p})$ that is conjugate to $y$. We included $\sqrt{-g(x)}$ so that the integral $\int_{y}$ is invariant under linearized diffeomorphism.  [^2] Diffeomorphism covariance to non-linear order in the diffeomorphism transform parameter $\xi_{\nu}$ requires special care because the midpoint $x$ used in defining does not transform covariantly. While several covariant prescriptions exist, notably via introducing horizontal lift to $e^{-i{\cal P}\cdot y}$ PhysRevD.37.2901 ; Fonarev:1993ht ; Antonsen:1997dc ; Liu2018 ; Hayata2020 , we have confirmed that these refinements yield identical results in our linear-response analysis and are therefore omitted for simplicity.
@@ 4.2 EM-response
$\delta^{{\rm EM}}_{L}\,W({\cal P},Q)=-A\cdot\partial_{{\cal P}}W_{0}({\cal P})=A\cdot\partial_{{\cal P}}S_{rr}({\cal P})$ (36)
@@
$\,\frac{i\,\delta_{{\cal P}_{+}}}{-P^{2}_{-}-m^{2}-i\epsilon p^{0}_{-}}=\frac{i\,\delta_{{\cal P}_{+}}}{-P^{2}_{+}-m^{2}+(P^{2}_{+}-P^{2}_{-})-i\epsilon p^{0}_{-}}=\frac{i\,\delta_{{\cal P}_{+}}}{2P\cdot Q-i\epsilon p^{0}_{-}}\,.$ (39)
@@ 5.1 Set-up
In fact, we can employ the aforementioned freedom to identify the non-analytic component, thereby making the matching more transparent. The numerator in the expression for $(\delta f,{\bm{g}})$ is proportional to the spatial gradient because of the P.B. structure in the kinetic equation. On the other hand, the numerator of $(\delta{\cal V},\delta A)_{\rm n.a.}$ (54) is proportional to $(Q\cdot u)f^{\prime}_{0}$, originating from a change in $p^{0}$ by the amount $\omega$ due to the temporal variation of external fields. Despite this difference, both expressions share the same non-analytic behavior upon using eq. (56) (i.e., upon integrating over $\int_{p^{0}}$).
@@ 5.2 Vector Wigner function
$\delta{\cal V}^{\mu}_{G}={\cal V}^{\mu}_{{\rm con}}[\delta f_{G}]+\textrm{analytic piece}\,,\qquad\delta\epsilon_{\rm G}=-\frac{1}{2}\,\epsilon_{\bm{p}}\,h_{\mu\nu}v^{\mu}v^{\nu}$ (60)
@@
Reversing the logic, this implies that diffeomorphism covariance also fixes the form of $\delta\epsilon_{G}$.
@@ 5.3 Axial Wigner function
We first evaluate the response to the EM perturbation ${\cal A}_{{\rm EM}}$ and metric perturbation ${\cal A}_{G}$. Using eq. (52) and eq. (43), we find
${\cal A}^{\mu}_{{\rm EM}}$ $=2\int_{p^{0}}\tilde{F}^{\mu\nu}\left[-v_{\nu}\frac{\bm{v}\cdot\bm{q}}{v\cdot{\cal Q}}f_{0}^{\prime}(p^{0})+v_{\nu}f^{\prime}_{0}(p^{0})-f(p^{0})\partial^{{\cal P}}_{\nu}\right]\delta^{+}_{{\cal P}}{}$
@@
${\cal A}^{\mu}_{G}=\frac{\bm{v}\cdot\bm{q}}{v\cdot Q}\,f_{0}^{\prime}\,(\frac{i}{2}\epsilon^{\mu\sigma\alpha\beta}v_{\alpha}Q_{\beta}\,h_{\sigma\nu}v^{\nu})-f^{\prime}_{0}\,(\frac{i}{2}\epsilon^{\mu\sigma\alpha\beta}v_{\alpha}Q_{\beta}h_{\sigma\nu}v^{\nu}_{\perp})\,\,.$ (71)
@@
${\bm{{\cal A}}}_{{\rm EM}}$ $={\bm{{\cal A}}}_{{\rm con}}[{\bm{g}}_{{\rm EM}},f_{\rm cov}]-f^{\prime}_{0}\,\frac{v^{2}}{\epsilon_{\bm{p}}}\bm{B}_{\perp}-\frac{f_{0}}{\epsilon_{\bm{p}}^{2}}(\bm{v}\times\bm{E})\,$, (72a)
$\bm{{\cal A}}_{\rm G}$ $={\bm{{\cal A}}}_{{\rm con}}[{\bm{g}}_{\rm G},f_{G}]+\left[-v^{2}\bm{B}^{h}_{\perp}+(\bm{v}\times\bm{E}^{h})+\bm{v}\times\nabla(\bm{v}\cdot\bm{A}^{h})\right]\,f^{\prime}_{0}\,$, (72b)
${\cal A}^{0}_{{\rm EM}}$ $={\cal A}^{0}_{\rm con}[{\bm{g}}_{{\rm EM}}]-\frac{\bm{v}\cdot\bm{B}}{\epsilon^{2}_{\bm{p}}}\,f_{0}\,$, (72c)
${\cal A}^{0}_{G}$ $={\cal A}^{0}_{\rm con}[{\bm{g}}_{G}]\,$, (72d)
@@
Importantly, the specific values of coefficients ($c$s) in ${\cal A}_{\rm con}$ (22b) do NOT depend on the types of perturbations and are given by
$c_{0}=c_{1}=1\,,\qquad c_{2}=(1-\bm{v}^{2})\qquad c_{3}=\frac{1}{\epsilon_{\bm{p}}}\,.$ (75)
@@
$\bm{v}\times{\bm{\partial}}f_{\rm cov}=\frac{\bm{v}\cdot q}{v\cdot Q}(\bm{v}\times\bm{E}-v^{2}\bm{B}_{\perp})f^{\prime}_{0}+v^{2}\bm{B}_{\perp}f^{\prime}_{0}\,\,.{}$
@@
$\bm{v}\times{\bm{\partial}}(\bm{E}\cdot\bm{v})=(\bm{v}\cdot{\bm{\partial}})(\bm{v}\times\bm{E}-v^{2}\bm{B}_{\perp})+v^{2}(v\cdot\partial)\bm{B}_{\perp}\,.{}$
@@
In addition to ${\cal A}_{\rm con}$, which is related to the dynamics of the distribution, there is also an analytic component that is interesting in its own right. For example, in ${\bm{{\cal A}}}_{{\rm EM}}$ (72a), we observe a term $\bm{v}\times\bm{E}$. This term manifests the spin Hall effect (SHE) (see Refs. Hattori:2019ahi ; Liu2021 for discussion in the context of QCD-like plasma, and ref. Yamamoto:2023okm for its application in the neutrino transport in core-collapse supernovae). Similarly, an analogous term appears in ${\bm{{\cal A}}}_{G}$ (72b), namely $\bm{v}\times\bm{E}^{h}$. In the homogeneous limit, one finds that $\bm{E}^{h}_{j}=-(\partial_{t}h_{ij}v^{j})/2$. In hydrodynamic limits, the response to the fluid shear stress $\sigma_{ij}$ can be mimicked by activating $\partial_{t}h_{ij}v^{j}$. In this sense, the shear-induced polarization Fu:2021pok ; Becattini_2021 can be heuristically thought of as an analogue of the SHE.
Since ${\cal A}_{G}$ is invariant under linear diffeomorphism, we now show that the expression for ${\cal A}_{G}$ appears simpler when expressed in a diffeomorphism-covariant form. In particular, the analytic component should vanish at first order in the gradient, as the lowest-order diffeomorphism-covariant terms are the Riemann curvature tensors $R^{\mu\nu}_{\ \rho\alpha}\sim\mathcal{O}(\partial^{2})$.
@@
${\cal A}^{i}_{G}=\frac{i}{v\cdot Q}f_{0}^{\prime}\left[\frac{1}{2}\left(\tilde{R}^{0i}_{\ \ 0l}-\epsilon^{im}_{\,\,\,\,\,\,n}v_{m}R^{0n}_{\ \ 0l}\right)v^{l}\right],\quad{\cal A}^{0}_{\rm G}=\bm{{\cal A}}_{\rm G}\cdot\bm{v}.$ (79)

### v2

@@ 1 Introduction
In response, significant developments have been made in formulating the SKT for relativistic plasma Hattori:2019ahi ; Weickgenannt:2019dks ; Gao:2019znl ; Weickgenannt:2020aaf ; Weickgenannt:2021cuo ; Hayata2020 ; Sheng:2021kfc ; Yang:2020hri (see ref. Hidaka:2022dmn for a review). A key predecessor to SKT is chiral kinetic theory, which describes the distribution of axial density/helicity Son:2012wh ; Stephanov:2012ki ; Son:2012zy ; Chen:2012ca ; Gao:2012ix ; Chen:2013iga ; Manuel:2013zaa ; Manuel:2014dza ; Chen:2014cla ; Chen:2015gta ; Hidaka:2016yjf ; Huang:2018wdl ; Gao:2018wmr ; Ma:2022ins . Most of the references mentioned above, particularly those from recent years, derive SKT using what we shall refer to as the ”top-down” approach. This approach begins with the microscopic Schwinger-Dyson equation for the Wigner function of the Fermionic field and then extracts the relevant equations for its components.
@@ 2.2.1 Dynamical variables
Because of the first inequality in eq. (2), we shall concentrate on the collisionless regime and neglect the self-energy corrections, which are of the order of $m_{D}$. The second inequality ensures that a gradient expansion exists in powers of $q/T_{{\rm eff}}$. Note that this gradient expansion parameter differs from that in hydrodynamics, which is given by $q/\gamma_{R}$. We focus on an effective description based on kinetic theory outside the hydrodynamic regime, consistent with our set-up that the background distribution $f_{0}$ need not be in local thermal equilibrium.
Eq. (2) assumes the separation of scales: $m_{D},\gamma_{R}\ll T_{{\rm eff}}$, which is true for a weakly coupled plasma with coupling constant $g_{C}<1$ where $m_{D}\sim g_{C}T_{{\rm eff}}$ and $\gamma_{R}\sim g^{2}_{C}T_{\rm eff}$.
@@ 2.2.2 Poisson bracket and equation of motion
We assume that $\bm{g}$} has no background value, and consequently, neither does $\bm{\zeta}$}. In the absence of $\bm{\zeta}$}, the evolution of $\bm{g}$} is determined by $\epsilon$ in the same way as $f$. The presence of $\bm{\zeta}$ introduces the coupling between $\bm{g}$} and $f$ in their evolution. The term $\bm{g}\times\bm{\zeta}$ originates from the last term in eq. (6b) which is a manifestation of $SU(2)$ algebra of spin $1/2$ particle. This non-linear term implies that if $\bm{\zeta}\propto\bm{B}$, $\bm{g}$ will evolve in time unless it aligns with the magnetic field $\bm{B}$ even if other terms like $\{{\bm{g}},\epsilon\}$ vanish.
Summarizing, we adopt a ”bottom-up” viewpoint by proposing that dynamical variables of SKT, $(f,{\bm{g}})$, obey E.o.Ms which are expressible in terms of spin-dependent P.B. Those equations have to be supplemented by the expression for the particle/spin energy $(\epsilon,{\bm{\zeta}})$. In general, $(\delta\epsilon,{\bm{\zeta}})$ are functionals of $(f,{\bm{g}})$ and external fields, similar to the framework of Landau Fermi liquid theory. The specification of those dependencies will be fixed by matching to the microscopic calculations; further details can be found in sec. 5.
@@ 2.2.3 Schwinger-Keldysh action
While one might think that $\Phi_{a}$ merely acts as the Lagrangian multiplier that ensures only solutions to E.o.M. contribute to the path integral  [^1] If we were using Martin–Siggia–Rose (MSR) method Martin:1973zz , we would place $i\phi_{a}$ where $\phi_{a}$ appears. However, the second condition in (9) requires that we shall use $\Phi_{a}$ without the prefactor $i$. , we emphasize that both $\Phi_{a}$ and its r-variable counterpart $\Phi$ possess much richer physical content in this formulation. This is reflected through the fact that various symmetry transformations can be implemented by the linear shift of $\Phi_{r/a}$. Particularly, we are interested in $U(1)$ gauge transformation $A_{\mu}\to A_{\mu}+\partial_{\mu}\alpha(x)$ and space-time translation parametrized by $b_{\mu}$. We further assume that $f$ describes the particles that carry the same $U(1)$ charge. By construction, $U(1)$ transformation is represented by (see also Ref Delacretaz:2025ifh )
@@ 2.4.2 Constitutive relation for axial Wigner function
We have projected $\bm{g}$} in the direction longitudinal and transverse to the three velocity $\bm{v}$, i.e., ${\bm{g}}_{\parallel}\equiv\hat{\bm{v}}(\hat{\bm{v}}\cdot{\bm{g}}),{\bm{g}}_{\perp}\equiv{\bm{g}}-{\bm{g}}_{\parallel}$. The coefficients $c_{0,1,2,3}$ are function of $\epsilon_{\bm{p}}$. We shall match eq. (22) with the field theoretical calculations later.
@@ 3 Linear response from kinetic theory
We shall match the kinetic theory description of the medium’s response to external disturbance with the field theory. In kinetic theory, the response is characterized by the deviations from the background distribution, i.e. $(\delta f,{\bm{g}})$ where $f=f_{0}+\delta f$ and $\bm{g}$}. Likewise, the presence of external fields also induces the change in energy, i.e., $\delta\epsilon$ in $\epsilon=\epsilon_{\bm{p}}+\delta\epsilon$ and $\bm{\zeta}$}. In our set-up, ${\bm{g}},\bm{\zeta}$ has no background value so we do not put $\delta$ in those symbols for brevity.
To analyze linear response, we set up power counting to keep track of the external perturbation by introducing a small parameter $\delta$. The background value of the distribution function $f_{0},{\cal V}^{\mu}_{0}\sim{\cal O}(\delta^{0})$ while $(\delta\epsilon,\bm{\zeta},{\bm{g}})\sim{\cal O}(\delta)$. For consistency, the function that parametrizes the symmetry transformation is also ${\cal O}(\delta)$. While both $\delta$ and gradient are counted as small, they are independent. For example, among terms which are linear in $\delta$, $\delta f,\delta\epsilon\sim{\cal O}(\partial^{0})$ while ${\bm{g}},\bm{\zeta}\sim{\cal O}(\partial)$.
@@
$\bm{g}(Q)$ $=\frac{\bm{v}\cdot\bm{q}}{v\cdot Q}\,f_{0}^{\prime}\,\bm{\zeta}(Q)\,$, (25b)
where from now on we assume that $f_{0}$ is a function of of $\epsilon_{\bm{p}}$ and $f^{\prime}_{0}$ denotes the derivative with respect to $\epsilon_{\bm{p}}$. We note from eqs. (25a),(25b) that $(\delta f,g)$ are non-analytic function in frequency and wave-vector $\omega,\bm{q}$, the point we shall return to in sec. 5. Note also that $(\delta f,{\bm{g}})$ in eq. (25) are proportional to $f^{\prime}_{0}$, which in the low temperature limit only receives contributions from excitations in the vicinity of the Fermi surface. The general lesson here is that the present EFT describes hard particles that contribute most to the deviation from the background value. The phase-space volume occupied by those excitations can be significantly smaller than that occupied by the particles in the bulk, indicating a reduction in the density of d.o.f.s governing the system’s dynamics.
@@ 4.1 Set-up
and $\sigma^{bc}=\frac{i}{2}[\gamma^{b},\gamma^{c}]_{-}$ are the generator of Lorentz transformation. The spin-connection term ensures the covariance of $W(x,{\cal P})$ under local Lorentz transformations (LLTs). The Wigner (Fourier) transform has been performed so that $W$ depends on both $x$ and the off-shell four-momentum ${\cal P}_{\mu}=(p_{0},\bm{p})$ that is conjugate to $y$. We included $\sqrt{-g(x)}$ so that the integral $\int_{y}$ is invariant under linearized diffeomorphism.  [^2] Diffeomorphism covariance to non-linear order in the diffeomorphism transform parameter $\xi_{\nu}$ requires special care because the midpoint $x$ used in defining does not transform covariantly. While several covariant prescriptions exist, notably via introducing horizontal lift to $e^{-i{\cal P}\cdot y}$ PhysRevD.37.2901 ; Fonarev:1993ht ; Antonsen:1997dc ; Liu2018 ; Hayata2020 , we have confirmed that these refinements yield identical results in our linear-response analysis and are therefore omitted for simplicity.
@@ 4.2 EM-response
$\delta_{L}\,W({\cal P},Q)=-A\cdot\partial_{{\cal P}}W_{0}({\cal P})=A\cdot\partial_{{\cal P}}S_{rr}({\cal P})$ (36)
@@
$\,\frac{i\,\delta_{{\cal P}_{+}}}{-{\cal P}^{2}_{-}-m^{2}-i\epsilon p^{0}_{-}}=\frac{i\,\delta_{{\cal P}_{+}}}{-{\cal P}^{2}_{+}-m^{2}+({\cal P}^{2}_{+}-{\cal P}^{2}_{-})-i\epsilon p^{0}_{-}}=\frac{i\,\delta_{{\cal P}_{+}}}{2{\cal P}\cdot Q-i\epsilon p^{0}_{-}}\,.$ (39)
@@ 5.1 Set-up
In fact, we can employ the aforementioned freedom to identify the non-analytic component, thereby making the matching more transparent. The numerator in the expression for $(\delta f,{\bm{g}})$ is proportional to the spatial gradient because of the P.B. structure in the kinetic equation. On the other hand, the numerator of $(\delta{\cal V},{\cal A})_{\rm n.a.}$ (54) is proportional to $(Q\cdot u)f^{\prime}_{0}$, originating from a change in $p^{0}$ by the amount $\omega$ due to the temporal variation of external fields. Despite this difference, both expressions share the same non-analytic behavior upon using eq. (56) (i.e., upon integrating over $\int_{p^{0}}$).
@@ 5.2 Vector Wigner function
$\delta{\cal V}^{\mu}_{\rm G}={\cal V}^{\mu}_{{\rm con}}[\delta f_{\rm G}]+\textrm{analytic piece}\,,\qquad\delta\epsilon_{\rm G}=-\frac{1}{2}\,\epsilon_{\bm{p}}\,h_{\mu\nu}v^{\mu}v^{\nu}$ (60)
@@
Reversing the logic, this implies that diffeomorphism covariance also fixes the form of $\delta\epsilon_{\rm G}$.
@@ 5.3 Axial Wigner function
We first evaluate the response to the EM perturbation ${\cal A}_{{\rm EM}}$ and metric perturbation ${\cal A}_{\rm G}$. Using eq. (52) and eq. (43), we find
${\cal A}^{\mu}_{{\rm EM}}$ $=2\int_{p^{0}}\tilde{F}^{\mu\nu}\left[-v_{\nu}\frac{\bm{v}\cdot\bm{q}}{v\cdot Q}f_{0}^{\prime}(p^{0})+v_{\nu}f^{\prime}_{0}(p^{0})-f(p^{0})\partial^{{\cal P}}_{\nu}\right]\delta^{+}_{{\cal P}}{}$
@@
${\cal A}^{\mu}_{\rm G}=\frac{\bm{v}\cdot\bm{q}}{v\cdot Q}\,f_{0}^{\prime}\,(\frac{i}{2}\epsilon^{\mu\sigma\alpha\beta}v_{\alpha}Q_{\beta}\,h_{\sigma\nu}v^{\nu})-f^{\prime}_{0}\,(\frac{i}{2}\epsilon^{\mu\sigma\alpha\beta}v_{\alpha}Q_{\beta}h_{\sigma\nu}(v^{\nu}-u^{\mu}))\,\,.$ (71)
@@
${\bm{{\cal A}}}_{{\rm EM}}$ $={\bm{{\cal A}}}_{{\rm con}}[{\bm{g}}_{{\rm EM}},f_{\rm cov}]-f^{\prime}_{0}\,\frac{\bm{v}^{2}}{\epsilon_{\bm{p}}}\bm{B}_{\perp}-\frac{f_{0}}{\epsilon_{\bm{p}}^{2}}(\bm{v}\times\bm{E})\,$, (72a)
$\bm{{\cal A}}_{\rm G}$ $={\bm{{\cal A}}}_{{\rm con}}[{\bm{g}}_{\rm G},f_{\rm G}]+\left[-\bm{v}^{2}\bm{B}^{h}_{\perp}+(\bm{v}\times\bm{E}^{h})+\bm{v}\times\nabla(\bm{v}\cdot\bm{A}^{h})-\bm{B}^{h}\right]\,f^{\prime}_{0}+\Delta{\bm{{\cal A}}}_{\rm G}\,$, (72b)
${\cal A}^{0}_{{\rm EM}}$ $={\cal A}^{0}_{\rm con}[{\bm{g}}_{{\rm EM}}]+\frac{f_{0}}{\epsilon^{2}_{\bm{p}}}(\bm{v}\cdot\bm{B})\,$, (72c)
${\cal A}^{0}_{\rm G}$ $={\cal A}^{0}_{\rm con}[{\bm{g}}_{\rm G}]-f_{0}^{\prime}(\bm{v}\cdot\bm{B}^{h})+\Delta{\cal A}^{0}_{\rm G}\,$, (72d)
@@
We also defined $\Delta{\cal A}^{\mu}_{\rm G}\equiv f^{\prime}_{0}(\frac{i}{2}\epsilon^{\mu\sigma\alpha\beta}v_{\alpha}Q_{\beta}h_{\sigma 0})$, coming from the term proportional to $u^{\mu}$ in eq. (71). Importantly, the specific values of coefficients ($c$s) in ${\cal A}_{\rm con}$ (22b) do NOT depend on the types of perturbations and are universally given by
$c_{0}=c_{1}=1\,,\qquad c_{2}=(1-\bm{v}^{2})\,,\qquad c_{3}=\frac{1}{\epsilon_{\bm{p}}}\,.$ (75)
@@
$\bm{v}\times{\bm{\partial}}f_{\rm cov}=\frac{\bm{v}\cdot q}{v\cdot Q}(\bm{v}\times\bm{E}-\bm{v}^{2}\bm{B}_{\perp})f^{\prime}_{0}+\bm{v}^{2}\bm{B}_{\perp}f^{\prime}_{0}\,\,.{}$
@@
$\bm{v}\times{\bm{\partial}}(\bm{E}\cdot\bm{v})=(\bm{v}\cdot{\bm{\partial}})(\bm{v}\times\bm{E}-\bm{v}^{2}\bm{B}_{\perp})+\bm{v}^{2}(v\cdot\partial)\bm{B}_{\perp}\,.{}$
@@
In addition to ${\cal A}_{\rm con}$, which is related to the dynamics of the distribution, there is also an analytic component that is interesting in its own right. For example, in ${\bm{{\cal A}}}_{{\rm EM}}$ (72a), we observe a term $\bm{v}\times\bm{E}$. This term manifests the spin Hall effect (SHE) (see Refs. Hattori:2019ahi ; Liu2021 for discussion in the context of QCD-like plasma, and ref. Yamamoto:2023okm for its application in the neutrino transport in core-collapse supernovae). Similarly, an analogous term appears in ${\bm{{\cal A}}}_{\rm G}$ (72b), namely $\bm{v}\times\bm{E}^{h}$. In the homogeneous limit, one finds that $\bm{E}^{h}_{j}=-(\partial_{t}h_{ij}v^{j})/2$. In hydrodynamic limits, the response to the fluid shear stress $\sigma_{ij}$ can be mimicked by activating $\partial_{t}h_{ij}v^{j}$. In this sense, the shear-induced polarization Fu:2021pok ; Becattini_2021 can be heuristically thought of as an analogue of the SHE.
Since ${\cal A}_{\rm G}$ is invariant under linear diffeomorphism, we now show that the expression for ${\cal A}_{\rm G}$ appears simpler when expressed in a diffeomorphism-covariant form. In particular, the analytic component should vanish at first order in the gradient, as the lowest-order diffeomorphism-covariant terms are the Riemann curvature tensors $R^{\mu\nu}_{\ \ \rho\alpha}\sim\mathcal{O}(\partial^{2})$.
@@
${\cal A}^{i}_{\rm G}=\frac{i}{v\cdot Q}f_{0}^{\prime}\left[\frac{1}{2}\left(\tilde{R}^{0i}_{\ \ 0l}-\epsilon^{im}_{\,\,\,\,\,\,n}v_{m}R^{0n}_{\ \ 0l}\right)v^{l}\right],\quad{\cal A}^{0}_{\rm G}=\bm{{\cal A}}_{\rm G}\cdot\bm{v}.$ (79)